# Fairness under Competition

### v1

@@ 1 Introduction
Algorithmic decision-making systems are not immune to human prejudices. This has been demonstrated by ample empirical evidence: For example, the use of algorithmic decision-making in determining loan approval and interest rates has led to minority applicants facing higher loan rejection rates and higher interest rates than non-minority applicants (Bartlett et al., 2022; Fuster et al., 2022), and credit card issuers offering lower credit limits for women than men (Gupta, 2019; Telford, 2019). Additional contexts in which such algorithmic bias has been documented include job recruitment (Doleac and Hansen, 2016), insurance premiums and payouts (Angwin et al., 2017), college admissions (Baker and Hawn, 2022; Gándara et al., 2024), and more (O’Neil, 2016).
@@
The regulator’s goal in imposing fairness adjustments to classifiers is thus to enhance the welfare of historically disadvantaged populations. One principled approach to fairness adjustment involves specifying a utility function, and requiring the classifier to ensure equal welfare across groups. Prominent fairness criteria such as Demographic Parity (DP) (Agarwal et al., 2018; Dwork et al., 2012) and Equal Opportunity (EO) (Hardt et al., 2016) can be interpreted as special cases within this welfare-based framework. For instance, DP corresponds to a utility specification where each approved applicant gains a utility of one and zero otherwise. EO, correcting for some criticisms about DP, does the same but only for the restricted set of deserving applicants.
@@
We begin in Section 2 by developing a model of ecosystem fairness. We suppose there are multiple lenders who make loan offers to borrowers; however, the model applies analogously to job candidates and employers, insurance buyers and insurers, or student applicants and colleges. We then use the welfare-based framework to define the notion of fairness under competition. In words, an ecosystem satisfies fairness under competition if the welfare of different groups of borrowers is equalized when the lenders use their classifiers to make loan offers. This definition is general, but we then specialize the definition to be the competition analogue to EO—namely, Equal Opportunity under Competition (EOC). The level of EOC here is a measure of how far the ecosystem is from satisfying EOC, where a lower level implies higher ecosystem fairness.
In Section 3 we turn to study EOC. Our main results are that, even when the individual classifiers satisfy EO, the ecosystem may be far from EOC. We quantify how far, based on model primitives, and provide worst-case bounds. In particular, we identify two distinct forces under which EO classifiers do not satisfy EOC. The first force, analyzed in Section 3.1, arises from differences in the correlation between the classifiers on the protected groups. The second force, analyzed in Section 3.2, arises when the lenders’ pools of potential borrowers are overlapping but not identical. We quantify the extent to which each force decreases ecosystem fairness.
@@
There are two lenders and two groups of borrowers. Each lender serves all applicants of group 1 but only to a subset of applicants of group 2, perhaps because the lenders have lower market penetration in the latter population. For served applicants, lenders use an EO classifier to predict loan repayment and offer loans to individuals predicted to repay.
@@
Finally, in order to explore the empirical prevalence of our theoretical results, in Section 5 we describe several experiments we ran on Lending Club loan data. In these experiments we compared the level of EOC before and after fairness adjustments under several variations. Our results indicate that, surprisingly, imposing fairness adjustments often leads to a higher level of EOC, and so lower ecosystem fairness.
@@
Our paper is part of a large and growing literature on fairness in machine learning (see, for example, Chouldechova and Roth, 2018; Mehrabi et al., 2021; Pessach and Shmueli, 2022; Barocas et al., 2023, among others). The literature includes a plethora of demonstrations of bias in machine learning, a large number of fairness metrics, and many algorithms for bias mitigation (the last of which is surveyed extensively by Hort et al., 2024).
Within this large literature, our paper fits within a subset of papers that question whether fair machine learning algorithms actually achieve fairness (see Ruggieri et al., 2023, for a survey). The paper perhaps most closely related to ours in this vein is that of Liu et al. (2018), who show that imposing fairness criteria may have adverse long-term effects. This can be seen as illustrating the failure of imposed fairness criteria due to temporal or sequential effects. Our paper, in contrast, can be seen as illustrating the failure due to competition or parallel effects. To the best of our knowledge, our paper is the first to examine fairness in a setting with multiple competing classifiers. The only somewhat related paper of which we are aware is that of Ustun et al. (2019), who compare the benefits to protected groups when there is one classifier with imposed fairness criteria, as opposed to many group-tailored individual classifiers. Although the latter case involves multiple classifiers, individuals are not classified by more than one, and so there is no competition between the classifiers.
@@ 2 Model
There are two types of players, borrowers and lenders. Each borrower is characterized by a triplet $(x,a,y)\in X\times A\times Y$, where $x$ is a vector of observable features, $a$ is an observable protected attribute, and $y$ is an unobservable outcome variable. For simplicity, assume $A=Y=\{0,1\}$. There is an underlying distribution $\mathcal{D}$ over $X\times A\times Y$. We slightly abuse notation by using $X$, $A$, and $Y$ as the sets of features, protected attributes, and outcome variables, and also as random variables with joint distribution $\mathcal{D}$ over the respective sets. Throughout, we think of each $(x,a,y)\in X\times A\times Y$ as an individual borrower.
There is a set $L$ of lenders. Each lender $\ell\in L$ decides whether or not to offer a loan to each borrower based on observables $(x,a)$. To do so the lender uses a potentially randomized classifier $c_{\ell}:X\times A\mapsto\{0,1\}$, where $c_{\ell}(x,a)=1$ if the lender offers borrower $(x,a)$ a loan, and $c_{\ell}(x,a)=0$ otherwise. The unobservable outcome variable $Y$ signifies whether or not the borrower repaid the loan, and lenders prefer to offer loans only to borrowers with $y=1$. Hence, perfect classifiers are ones where $c_{\ell}(x,a)=y$ for each borrower $(x,a,y)$. In general, however, firms’ classifiers are not perfect. For a given classifier $c_{\ell}$, denote its false-negative rate by $\beta_{\ell}=\Pr\left[c_{\ell}(X,A)=0|Y=1\right]$ and its false-positive rate by $\alpha_{\ell}=\Pr\left[c_{\ell}(X,A)=1|Y=0\right]$. Conversely, $1-\beta_{\ell}$ and $1-\alpha_{\ell}$ are the classifier’s true-negative and true-positive rates, respectively.
Borrowers’ utilities are described by a function $v:X\times A\times Y\times\{0,1\}^{\left|{L}\right|}\mapsto{\mathbb{R}}$, where $v\left(x,a,y,o_{1},\ldots,o_{\left|{L}\right|}\right)$ is the utility of borrower $(x,a,y)$ when he receives offers from all lenders $\ell$ with $o_{\ell}=1$. For simplicity, we will assume henceforth that borrowers’ utilities depend only on the number of offers they received, and so redefine the utility functions as $v:X\times A\times Y\times\mathbb{N}\mapsto{\mathbb{R}}$, where $\mathbb{N}$ is the set of non-negative integers. Thus, $v(x,a,y,r)$ is the utility of a borrower $(x,a,y)$ when he receives $r$ loan offers.
@@
We begin with the general welfare-equalizing (WE) notion of fairness of Ben-Porat et al. (2021) and the special case of EO. EO is derived from WE by fixing utilities $v(x,y,a,r)=1$ if $y=r=1$, and $v(x,y,a,r)=0$ otherwise.
  • WE: For a given utility function $v$ of the borrowers and classifier $c_{\ell}$ of a lender, denote by $W_{v,c_{\ell}}(a)=\mathrm{E}\left[v(X,A,Y,1)\cdot c_{\ell}(X,A)|A=a\right]$. The $v$-WE level of a classifier $c_{\ell}$ is $\left|{W_{v,c_{\ell}}(0)-W_{v,c_{\ell}}(1)}\right|$, the difference in welfare between the two groups under classifier $c_{\ell}$.
  • EO: The EO level of $c_{\ell}$ is $\left|{\mathrm{E}\left[c_{\ell}(X,A)|Y=1,A=0\right]-\mathrm{E}\left[c_{\ell}(X,A)|Y=1,A=1\right]}\right|$, the difference in the fraction of offers made to each of the groups’ “deserving” borrowers. A classifier is EO if its EO level is 0.
@@
We now generalize the definitions of fairness to a setting with multiple competing lenders, defining the WE under competition (WEC) and EO under competition (EOC). Fix a utility function $v$ and classifiers $c=(c_{1},\ldots,c_{\left|{L}\right|})$.
  • WEC: For a given utility function $v$ of the borrowers and classifiers $c_{1}$ and $c_{2}$ of the lenders, denote by $W_{v,c}(a)=E\left[v(X,A,Y,R(X,A,Y))|A=a\right]$, where $R(x,a,y)=\sum_{\ell\in L}c_{\ell}(x,a)$ is the number offers given to borrower $(x,a,y)$. The $v$-WE level of classifiers $c$ is then $\left|{W_{v,c}(0)-W_{v,c}(1)}\right|$.
We now impose several restrictions for simplicity, but note that the technical appendix extends some results to the more general setting. First, we assume that there are only two lenders, and so that $L=\{1,2\}$. Second, we focus on the case in which borrowers do not care which or how many offers they obtain, as long as they obtain at least one. Thus, $v(x,a,y,1)=v(x,a,y,2)=1$, and $v(x,a,y,0)=0$. Finally, we focus on a particular notion of ecosystem fairness, namely, EOC. For this definition, denote by $d(x,a)=\Pr\left[R(x,a,y)\geq 1\right]$, the probability that at least one lender offers borrower $(x,a)$ a loan.
  • EOC: The EOC level of the classifiers is
    $\left|{\mathrm{E}\left[d(X,A)|Y=1,A=0\right]-\mathrm{E}\left[d(X,A)|Y=1,A=1\right]}\right|,$
    the difference between the two groups in the fraction of “deserving” borrowers who obtained at least one offer. Classifiers are EOC if their EOC level is 0.
@@ 3 Equal opportunity under competition
In this section we ask the following question: Given two EO classifiers, are they guaranteed to be EOC? We show that the answer is negative and identify two distinct forces that drive this. We also quantify the level of EOC, and bound the worst case. The first force, analyzed in Section 3.1, arises from differences in the correlation between the classifiers on the two protected groups. The second force, analyzed in Section 3.2, arises when each classifier serves only a subset of the borrowers.
@@ 3.1 Correlations between classifiers
Fix two EO classifiers $c_{1}$ and $c_{2}$ with false-negative rates $\beta_{1}$ and $\beta_{2}$. We define two Bernoulli random variables that capture the true positives of the classifiers on the two groups $a\in A$. Specifically, for each $a\in A$ and each $\ell\in\{1,2\}$, define the Bernoulli random variable
@@
Each $B_{\ell}^{a}$ is the output of classifier $c_{\ell}$ on random instances with $A=a$ and $Y=1$, and so $\mathrm{E}\left[B_{\ell}^{a}\right]=1-\beta_{\ell}$ is the true-positive rate of classifier $c_{\ell}$. Denote by $\sigma_{\ell}=\sqrt{\beta_{\ell}(1-\beta_{\ell})}$ the standard deviation of $B_{\ell}^{0}$, and note that, because $c_{\ell}$ is EO, $\sigma_{\ell}$ is also the standard deviation of $B_{\ell}^{1}$. Finally, for each $a$ let $\rho^{a}$ denote the Pearson correlation coefficient between $B_{1}^{a}$ and $B_{2}^{a}$. The Pearson correlation captures the extent to which true positives by one classifier are correlated with true positives by the other classifier.

### v2

@@ 1 Introduction
Algorithmic decision-making systems are not immune to human prejudices. This has been demonstrated by ample empirical evidence: For example, the use of algorithmic decision-making in determining loan approval and interest rates has led to minority applicants facing higher loan rejection rates and higher interest rates than non-minority applicants (Bartlett et al., 2022; Fuster et al., 2022), and credit card issuers offering lower credit limits for women than men (Gupta, 2019; Telford, 2019). Additional contexts in which such algorithmic bias has been documented are as diverse as job recruitment (Doleac and Hansen, 2016), insurance premiums and payouts (Angwin et al., 2017), college admissions (Baker and Hawn, 2022; Gándara et al., 2024), and more (O’Neil, 2016).
@@
The regulator’s goal in imposing fairness adjustments to classifiers is thus to enhance the welfare of historically disadvantaged populations. One principled approach to fairness adjustment involves specifying a utility function, and requiring the classifier to ensure equal welfare across groups. Prominent fairness criteria such as Demographic Parity (DP) (Agarwal et al., 2018; Dwork et al., 2012), as well as Equal Opportunity (EO) and Equalized Odds (ED) (Hardt et al., 2016) can be interpreted as special cases within this welfare-based framework. For instance, DP corresponds to a utility specification where each approved applicant gains a utility of one and zero otherwise. EO and ED, correcting for some criticisms about DP, does the same but only for the restricted set of deserving applicants.
@@
We begin in Section 2 by developing a model of ecosystem fairness. We suppose there are multiple lenders who make loan offers to borrowers; however, the model applies analogously to job candidates and employers, insurance buyers and insurers, or student applicants and colleges. For this model we define our main notion of fairness under competition, a competition analogue of EO that we call Equal Opportunity under Competition (EOC). The level of EOC here is a measure of how far the ecosystem is from satisfying EOC, where a lower level implies higher ecosystem fairness. We also define a second, welfare-based version of EOC. An ecosystem satisfies this second notion if the welfare of different groups of borrowers is equalized when the lenders use their classifiers to make loan offers.
In Section 3 we turn to study EOC. Our main results are that, even when the individual classifiers satisfy EO, the ecosystem may be far from EOC. We quantify how far, based on model primitives, and provide worst-case bounds. In particular, we identify two distinct forces under which EO classifiers do not satisfy EOC. The first force, analyzed in Section 3.1, arises from differences in the correlations between the classifiers on the protected groups. The second force, analyzed in Section 3.2, arises when the lenders’ pools of potential borrowers are overlapping but not identical. We quantify the extent to which each force decreases ecosystem fairness.
@@
There are two lenders and two groups of borrowers. Each lender serves all applicants of group 1 but only a subset of applicants of group 2, perhaps because the lenders have lower market penetration in the latter population. For served applicants, lenders use an EO classifier to predict loan repayment and offer loans to individuals predicted to repay.
@@
Section 3 focuses on EOC when there are two lenders. However, in Sections 3.1.1 and 3.1.2 we extend some results to the welfare-based notion of fairness and to an arbitrary number of lenders.
@@
Next, in order to explore the empirical prevalence of our theoretical results, in Section 5 we describe several experiments we ran on Lending Club loan data. In these experiments we compared the level of EOC before and after fairness adjustments under several variations. Our results indicate that, surprisingly, imposing fairness adjustments often leads to a higher level of EOC, and so lower ecosystem fairness.
Finally, in Appendix B we extend our results to other notions of fairness, namely, ED and DP.
@@
Our paper is part of a large and growing literature on fairness in machine learning (see, e.g., Chouldechova and Roth, 2018; Mehrabi et al., 2021; Pessach and Shmueli, 2022; Barocas et al., 2023, among others). The literature includes a plethora of demonstrations of bias in machine learning, a large number of fairness metrics, and many algorithms for bias mitigation (the last of which is surveyed extensively by Hort et al., 2024).
Within this large literature, our paper fits within a subset of papers that question whether fair machine learning algorithms actually achieve fairness (see Ruggieri et al., 2023, for a survey). One paper related to ours in this vein is that of Liu et al. (2018), who show that imposing fairness criteria may have adverse long-term effects. This can be seen as illustrating the failure of imposed fairness criteria due to temporal or sequential effects. Our paper, in contrast, can be seen as illustrating the failure due to competition or parallel effects.
Our research is more closely related to papers that examine fairness in settings with multiple classifiers. Two such papers, Bower et al. (2017) and Dwork and Ilvento (2019), also discuss the insight that fairness of individual classifiers does not imply fairness of a system composed of multiple classifiers. Our paper differs along multiple dimensions. First, in terms of motivation, these papers ask when a central platform (e.g., an ad platform) will be fair when handling a task that consists of several subtasks (advertisers on the platform), where each subtask is required to be fair. Our motivation is more about competing firms in a decentralized market, where the joint activity determines the user’s utility. More specifically, Dwork and Ilvento (2019) largely focus on individual fairness. They do have some extensions to group fairness, and their main results here are to show that there exist distributions for which individual fairness does not imply joint fairness. However, in their model the classifiers are assumed to always be independent, and they cannot capture the correlations between classifiers that lead to ecosystem unfairness. In addition, their notion of group fairness does not include EO. Bower et al. (2017) do focus on EO, but here the main difference is that classifiers are composed sequentially: one classifier makes a prediction or decision, the outcome is then passed on to the next classifier, and so on. In our setting, in contrast, the classifiers run in parallel. Finally, although both papers contain the insight that individual fairness does not suffice for joint fairness, in our paper we also identify the forces that lead to the joint unfairness, and quantify the extent of this unfairness as a function of correlations and overlaps.
Another somewhat related paper is that of Ustun et al. (2019), who compare the benefits to protected groups when there is one classifier with imposed fairness criteria, as opposed to many group-tailored individual classifiers. Although the latter case involves multiple classifiers, individuals are not classified by more than one, and so there is no competition between the classifiers.
@@ 2 Model
There are two types of players, borrowers and lenders. Each borrower is characterized by a triplet $(x,a,y)\in X\times A\times Y$, where $x$ is a vector of observable features, $a$ is an observable protected attribute (group membership), and $y$ is an unobservable outcome variable. For simplicity, assume $A=Y=\{0,1\}$. There is an underlying distribution $\mathcal{D}$ over $X\times A\times Y$. We slightly abuse notation by using $X$, AA, and $Y$ as the sets of features, protected attributes, and outcome variables, and also as random variables with joint distribution $\mathcal{D}$ over the respective sets. Throughout, we think of each $(x,a,y)\in X\times A\times Y$ as an individual borrower.
There is a set $L$ of lenders. Each lender $\ell\in L$ decides whether or not to offer a loan to each borrower based on observables $(x,a)$. To do so the lender uses a potentially randomized classifier $c_{\ell}:X\times A\mapsto\{0,1\}$, where $c_{\ell}(x,a)=1$ if the lender offers borrower $(x,a)$ a loan, and $c_{\ell}(x,a)=0$ otherwise. The unobservable outcome variable $Y$ signifies whether or not the borrower repays the loan, and we call borrowers with $y=1$ deserving borrowers. Naturally, lenders prefer to offer loans only to deserving borrowers. Hence, perfect classifiers are ones where $c_{\ell}(x,a)=y$ for each borrower $(x,a,y)$. In general, however, firms’ classifiers are not perfect. For a given classifier $c_{\ell}$, denote its false-negative rate by $\beta_{\ell}=\Pr\left[c_{\ell}(X,A)=0|Y=1\right]$ and its false-positive rate by $\alpha_{\ell}=\Pr\left[c_{\ell}(X,A)=1|Y=0\right]$. Conversely, $1-\beta_{\ell}$ and $1-\alpha_{\ell}$ are the classifier’s true-negative and true-positive rates, respectively.
Borrowers’ utilities are described by a function $v:X\times A\times Y\times\mathbb{N}\mapsto{\mathbb{R}}$, where $\mathbb{N}$ is the set of non-negative integers. $v(x,a,y,r)$ is the utility of a borrower $(x,a,y)$ when she receives $r$ offers. [^1] More generally, borrowers’ utilities could also depend on which lenders extended offers. Our definitions and results apply to this more general setting as well.
@@
We begin by stating two notions of fairness for the setting of a single classifier. [^2] These notions are the foci of the paper, but in Appendix B we extend our results to other notions. The first is the popular notion of EO (Hardt et al., 2016), which requires the classifier’s false-negative rates to be equal across groups. The second is based on the welfare-equalizing notion of Ben-Porat et al. (2021) that we call $v$-EO, and which requires the expected utilities of deserving borrowers to be equal across groups.
  • EO:
    The EO level of $c_{\ell}$ is $\left|{\mathrm{E}\left[c_{\ell}(X,A)|Y=1,A=0\right]-\mathrm{E}\left[c_{\ell}(X,A)|Y=1,A=1\right]}\right|$, the difference in the fraction of offers made to each of the groups’ deserving borrowers. A classifier is EO if its EO level is 0.
  • $v$-EO:
    For a given utility function $v$ of the borrowers and classifier $c_{\ell}$ of a lender, denote by $W_{v,c_{\ell}}(a)=\mathrm{E}\left[v(X,A,Y,c_{\ell}(X,A))|Y=1,A=a\right]$ the welfare of deserving borrowers from group $a$. The $v$-EO level of a classifier $c_{\ell}$ is $\left|{W_{v,c_{\ell}}(0)-W_{v,c_{\ell}}(1)}\right|$, the difference in welfare between the deserving borrowers of the two groups under classifier $c_{\ell}$.
@@
We now generalize the definitions of fairness to a setting with multiple competing lenders, EO under competition (EOC) and $v$-EO under competition ($v$-EOC) . Fix a utility function $v$ and classifiers $c=(c_{1},\ldots,c_{\left|{L}\right|})$, and denote by $R(x,a)=\sum_{\ell\in L}c_{\ell}(x,a)$ the number offers made to a borrower with observables $(x,a)$. Also, denote by $d(x,a)=\Pr\left[R(x,a)\geq 1\right]$, the probability that at least one lender offers a borrower with observables $(x,a)$ a loan.
  • EOC:
    The EOC level of classifiers $c$ is $\left|{\mathrm{E}\left[d(X,A)|Y=1,A=0\right]-\mathrm{E}\left[d(X,A)|Y=1,A=1\right]}\right|$, the difference between the two groups in the fraction of deserving borrowers who obtained at least one offer. Classifiers are EOC if their EOC level is 0.
  • $v$-EOC:
    For a given utility function $v$ of the borrowers and classifiers $c$ of the lenders, denote by $W_{v,c}(a)=E\left[v(X,A,Y,R(X,A))|Y=1,A=a\right]$. The $v$-EOC level of classifiers $c$ is then $\left|{W_{v,c}(0)-W_{v,c}(1)}\right|$.
In some of our analyses we will impose several assumptions. First, we sometimes focus on the case of two lenders. Second, we sometimes focus on the case in which borrowers do not care how many offers they obtain, as long as they obtain at least one. In this case, their preferences satisfy the following:
Assumption 1 (0-1 preferences)
$v(x,a,y,r)=1$ if $r\geq 1$ and $v(x,a,y,0)=0$, $\forall x,a,y$.
Observe that under Assumption 1, EO is equivalent to $v$-EO and EOC is equivalent to $v$-EOC. For more general utility functions EO and EOC may differ from $v$-EO and $v$-EOC.
@@ 3 Equal Opportunity under Competition
In this section we ask the following question: Given two EO classifiers, are they guaranteed to be EOC? We show that the answer is negative, and identify two distinct forces that drive this result. We also quantify the level of EOC, and bound the worst case. The first force, analyzed in Section 3.1, arises from differences in the correlation between the classifiers on the two protected groups. The second force, analyzed in Section 3.2, arises when each classifier serves only a subset of the borrowers.
@@ 3.1 Correlations between classifiers
Fix two EO classifiers $c_{1}$ and $c_{2}$ with false-negative rates $\beta_{1}$ and $\beta_{2}$. We define two Bernoulli random variables that capture the true positives of the classifiers on the two groups $a\in A$: For each $a\in A$ and each $\ell\in\{1,2\}$, let the Bernoulli random variable
@@
Each $B_{\ell}^{a}$ is the output of classifier $c_{\ell}$ on random instances with $A=a$ and $Y=1$, and so $\mathrm{E}\left[B_{\ell}^{a}\right]=1-\beta_{\ell}$ is the true-positive rate of classifier $c_{\ell}$. Denote by $\sigma_{\ell}=\sqrt{\beta_{\ell}(1-\beta_{\ell})}$ the standard deviation of $B_{\ell}^{0}$, and note that, because $c_{\ell}$ is EO, $\sigma_{\ell}$ is also the standard deviation of $B_{\ell}^{1}$. Finally, for each $a$ let $\rho^{a}$ denote the Pearson correlation coefficient between $B_{1}^{a}$ and $B_{2}^{a}$. The Pearson correlation captures the extent to which the true positives of the classifiers correlate with one another.